OpenStudy (anonymous):

simplify y^3/2 * y^3/2 * y^3/2

OpenStudy (anonymous):

\[x^a\cdot x^b=x^{a+b}\]

OpenStudy (anonymous):

so will it be y^9/8

OpenStudy (anonymous):

\[\frac{3}{2}+\frac{3}{2}+\frac{3}{2}\neq\frac{9}{8}\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw